Financing Instead of paying rent, UniVEND is consider purchasing several vending machines and operating them personally. To do this, Semih needs 500,000 TL in capital. Help Semih determine the best financing option for his business Part1 Option No. 1- Bank Loan Kar Bank has a special offer for SMEs, offering a 4-year loan at 1.1% What are the monthly payments if Semih borrows 500,000 TL? How much of a down payment must Semih pay if he doesnt want his monthly payments to exceed 10,000 TL monthly? Semih has the option to extend the term of the loan to 5 years, but the monthly interest rate will increase to 1.2%. How much does this change his monthly payments? (HINT: find the difference from question 1a) Semih also spoke with a friend of his who works at Güneş Bank and he received the following offer: monthly payments of 10,000 TL at a 1.05% monthly interest rate, followed by a lump sum payment of 60,000 TL at the end of the term. Payments are made at the end of the period. What is the total term of the loan? Give your answer in terms of months. interest rate, payments made at the beginning of the period b) If Semih wants to pay off the loan within 5 years (l.e. 60 months), how much should he increase the lump sum payment to be paid at the end of the term? Instead of increasing the lump sum payment, perhaps Semih will just borrow less. How many can he borrow if he wants to pay off the loan in 5 years? Part2 Option No. 2-Investor Capital Semih has found an investor who is willing to provide him with 500,000 TL in exchange for 30% ownership, which means he will take 30% of the annual profits. The investor generally seeks a 12% ROI for his investments. Lets say Semih anticipates that he will have future profits of 200,000 TL per year for the forseeable future. What is the Present Value of the investment for the Based on the investment amount, what is the total valuation of Semihs company? (Hint: if 30% 500,000 TL, then 100%·?) Semih has an ambitious growth plan and has put together the following profit projections see below). For 10 years, the investor will take 30% of the annual profits. After 10 years, the will exit the business by selling his shares back to Semih for 500,000 TL Report the NPV for the investor, using a 12% ROI Report the IRR for the investor vestor b) Profit Projections Year 100,000.00 125,000.00 156,000.00 195,000.00 244,000.00 305,000.00 381,000.00 77,000.00 596,000.00 745,000.00 10

Please only part 1)section 2 a.b and c. Also part 2) section 3 a,bic

please it is needed please solve
Answer #1

Option 1 - Section 2:

(a) If total borrowing is 500000 TL and repayment would involve monthly payouts of 10000 TL each and a final lump sum payout of 60000 TL, the total present value of all these repayment cash flows discounted at the applicable monthly interest rate of 1.05 % should equal the initial total borrowing of 500000 TL

Let the tenure of the repayments be N months

Therefore, 500000 = 10000 x (1/0.0105) x [1- {1/(1.0105)^(N)}] + 60000 / (1.0105)^(N)

500000 = 952381 x [1- {1/(1.0105)^(N)}] + 60000 / (1.0105)^(N)

Using trial and error/ EXCEL's Goal Seek Function to solve the equation we get: N = 65.041 ~ 65 months

(b) Let the lump sum payment be K and new repayment tenure is 60 months

500000 = 10000 x (1/0.0105) x [1-{1/(1.0105)^(60)}] + K / (1.0105)^(60)

K ~ 105770.98 TL

(c) Let the amount borowed be M TL

Therefore, M = 10000 x (1/0.0105) x [1-{1/(1.0105)^(60)}] + 60000 / (1.0105)^(60) = 475542.55 TL

Option 2 - Section 3:

(a) Required ROI = 12 % , Pertual Expected Profit = 200000 TL and Profit Share = 30 %

Investor's Perpetual Profit = 0.3 x 200000 = 60000 TL

Present Value of Investor's Investment = 60000 / 0.12 = 500000 TL
(b) 30% of the company can be bought for 500000 TL

Therefore, 100% (entire firm) of the firm can be bought for = Firm Value = 500000 / 0.3 ~ 1666666.67 TL